Evaluate the limits:
lim┬(x→0)⁡〖1/x^3 〗 ∫_0^x▒〖t^2/(t^4+1) dt〗
1
Expert's answer
2020-08-31T16:39:35-0400

If "0\leq t\leq x", then "1+t^4\leq 1+x^4" , hence

"1=\frac{1}{1+0^4}\geq\frac{1}{1+t^4}\geq\frac{1}{1+x^4}," hence

"\frac{1}{3}=\frac{1}{x^3}\intop_0^x t^2dt\geq\\
\frac{1}{x^3}\intop_0^x\frac{t^2}{1+t^4}dt\geq\frac{1}{x^3}\intop_0^x\frac{t^2}{1+x^4}dt=\frac{1}{3(1+x^4)}",

"\lim\limits_{x\rightarrow+0}\frac{1}{3}\geq\\
\lim\limits_{x\rightarrow+0}\frac{1}{x^3}\intop_0^x\frac{1}{1+t^4}dt\geq\\
\lim\limits_{x\rightarrow+0}\frac{1}{3(1+x^4)}=\frac{1}{3(1+0)}=\frac{1}{3},"

By The Squeeze Theorem

"\lim\limits_{x\rightarrow+0}\frac{1}{x^3}\intop_0^x\frac{t^2}{1+t^4}dt=\frac{1}{3}."

Similarly, if "x\leq t\leq 0", then "1+t^4\leq 1+x^4" ,

hence

"1=\frac{1}{1+0^4}\geq\frac{1}{1+t^4}\geq\frac{1}{1+x^4},"

hence for "x<0"

"\frac{1}{3}=\frac{1}{x^3}\intop_0^x t^2dt=-\frac{1}{x^3}\intop_x^0 t^2dt\geq\\
-\frac{1}{x^3}\intop_x^0\frac{t^2}{1+t^4}dt=\frac{1}{x^3}\intop_0^x\frac{t^2}{1+t^4}dt\geq-\frac{1}{x^3}\intop_x^0\frac{t^2}{1+x^4}dt=\\
\frac{1}{x^3}\intop_0^x\frac{t^2}{1+x^4}dt=\frac{1}{3(1+x^4)},"

hence

"\lim\limits_{x\rightarrow-0}\frac{1}{3}\geq\\
\lim\limits_{x\rightarrow-0}\frac{1}{x^3}\intop_0^x\frac{1}{1+t^4}dt\geq\\
\lim\limits_{x\rightarrow-0}\frac{1}{3(1+x^4)}=\frac{1}{3(1+0)}=\frac{1}{3},"

by The Squeeze Theorem

"\lim\limits_{x\rightarrow-0}\frac{1}{x^3}\intop_0^x\frac{t^2}{1+t^4}dt=\frac{1}{3},"

hence

"\lim\limits_{x\rightarrow0}\frac{1}{x^3}\intop_0^x\frac{1}{1+t^4}dt=\frac{1}{3}."

Answer: "\frac{1}{3}."